2.) 350 Individuals are surveyed regarding smoking habit and marital status.

Smoker Non-smoker Total Married 54 146 200 Divorced 38 62 100 Never Married 11 39 50 Total 103 247 350 a. If one subject is randomly selected, what is the probability of selecting a smoker? P(smoker) = b. If one subject is randomly selected, what is the probability of selecting someone who is both a smoker and divorced? P(smoker AND divorced) = c. If one subject is randomly selected, what is the probability of selecting someone who is either a smoker or is divorced? P(smoker OR divorced) = d. What is the probability that a selected individual is divorced, given that he is a smoker? P(divorced | smoker) =
Answer #1
smoker non smoker total
married 54 146 200
divorced 38 62 100
never married 11 39 50
total 103 247 350

a)  probability of selecting a smoker =103/350 =0.2943

b) probability of selecting someone who is both a smoker and divorced =38/350=0.1086

c)P(smoker OR divorced) =(103+100-38)/350 =165/350 =33/70 =0.4714

d)

P(divorced | smoker) =38/103 =0.3689